40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 13 09:15:15 2021, Paco, Update, LSC, MC unlocked? VAC_2021-09-13_09-32-45.png
    Reply  Mon Sep 13 15:14:36 2021, Anchal, Update, LSC, Xend Green laser injection mirrors M1 and M2 not responsive 
       Reply  Thu Sep 30 14:13:18 2021, Anchal, Update, LSC, HV supply to Xend Green laser injection mirrors M1 and M2 PZT restored 
Message ID: 16368     Entry time: Thu Sep 30 14:13:18 2021     In reply to: 16322
Author: Anchal 
Type: Update 
Category: LSC 
Subject: HV supply to Xend Green laser injection mirrors M1 and M2 PZT restored 

Late elog, original date Sep 15th

We found that the power switch of HV supply that powers the PZT drivers for M1 and M2 on Xend green laser injection alignment was tripped off. We could not find any log of someone doing it, it is a physical switch. Our only explanation is that this supply might have a solenoid mechansm to shut off during power glitches and it probably did so on Aug 23 (see 40m/16287). We were able to align the green laser using PZT again, however, the maximum power at green transmission from X arm cavity is now about half of what it used to be before the glitch. Maybe the seed laser on the X end died a little.

ELOG V3.1.3-